3. Suppose the marginal benefits (social and private) and costs (social and pri vate) per gallon of gasoline in Canada are modelled as follows to illustrate the negative externality of gasoline combustion: MSB- 12.80 0.42Q MPB- 12.80- 0.4Q MSC- MPC 1.25 +.02Q where Q is millions of gallons (a) State the equation that represents the market externality. Give the economic interpretation of this equation, using its specific numerical value(s) round off.) the efficient solution, and calculate the tax revenues generated to the (b) Find the efficient equilibrium, P* and Q, for this market. (Do not (c) Find the dollar value of a per-unit gasoline tax that would achieve government as result.
Answer #1

(a)

Externality is defined as

Marginal external benefit (MEB) = MSB - MPB

MEB = 12.8 - 0.42Q - (12.8 - 0.4Q) = 12.8 - 0.42Q - 12.8 + 0.4Q

MEB = -0.02Q

(b)

Efficient outcome is obtained by equating MSB and MSC.

12.8 - 0.42Q = 1.25 + 0.02Q

0.44Q = 11.55

QE = 26.25

PE = 1.25 + (0.02 x 26.25) = 1.25 + 0.525 = 1.775

(c)

When QE = 26.25, per-unit gasoline tax is absolute value of MEB.

Unit tax = 0.02 x 26.25 = $0.525

Tax revenue = QE x Unit tax = 26.25 x $0.525 = $13.78125